OpenStudy (anonymous):

Can someone please explain How do I solve this 2xe^x+x^2e^x= 0 e^3x+3xe^3x=0 2x^2 e^x^2- e^x^2 / x^2 =0

OpenStudy (anonymous):

just put all the numbers together

OpenStudy (anonymous):

33322222 that would be my answer

OpenStudy (anonymous):

what do have to do?

OpenStudy (anonymous):

put all the numbers you see together

OpenStudy (anonymous):

Solve the equations. There is three seperate equations so you can't put all the numbers together

OpenStudy (anonymous):

oh if it was me i would just put what i see

OpenStudy (anonymous):

\[2xe ^{x} +x^2 e^x =0\]

OpenStudy (anonymous):

so 0

OpenStudy (anonymous):

no. What ?

OpenStudy (campbell_st):

for the 1st question, take out the common factor of e^x then solve the quadratic \[e^{3x}(1 + 3x) = 0\] the problem is that \[e^{3x} \neq 0\] as you can't find a value of x that make e^(3x) zero. so then by taking the common factor you can find the value of x that makes 1 + x = 0 because if you substitute that value into the equation you would have e^{3x}*(0) = 0 hope this helps

OpenStudy (anonymous):

Thank you it does

OpenStudy (campbell_st):

is the 2nd question 1. \[2x^2e^{x^2} - \frac{e^{x^2}}{x^2} = 0\] or 2. \[\frac{2x^2 e^{x^2} - e^{x^2}}{x^2}=0\]

OpenStudy (campbell_st):

and for the 1st question take e^x as a common factor and you have \[e^x(2x + x^2) = 0\] solve the quadratic.... as the exponential will never equal zero

OpenStudy (anonymous):

The whole thing is being divided by x^2

OpenStudy (campbell_st):

ok.. so factor the numerator you get \[\frac{e^{x^2}(2x^2 -1)}{x^2}\] so now looking at the problem, the numerator can't be zero so just solve the quadratic \[2x^2 - 1 = 0\]

OpenStudy (anonymous):

Okay thank you so much